# Non-Stationary Lipschitz Bandits

### v1

@@ 4 Main results
Let $\mathcal{E}_{\mathrm{Lip}}(T,\tilde{L})$ denote the class of non-stationary Lipschitz bandit environments over the arm space $[0,1]$ with at most $\tilde{L}_{T}\leq L$ significant shifts. There exists a numerical constant $\underaccent{\bar}{c}_{1}>0$ such that for any algorithm $\pi$,
@@
To give intuition whether the optimal rate $\mathcal{O}(\tilde{L}^{1/3}T^{2/3})$ is achievable, consider an oracle algorithm $\pi_{\rm{oracle}}$ with access to the exact times of the significant shifts $\{\tau_{i}\}_{i=0}^{\tilde{L}_{T}}$. The optimal algorithm partitions the horizon into $\tilde{L}_{T}$ phases and, within each phase, runs a minimax-optimal bandit algorithm (e.g. UCB (Auer et al., 2002a)) using a discretization of $(\tau_{i+1}-\tau_{i})^{1/3}$ arms. This oracle algorithm suffers regret at most $c_{\rm{oracle}}(\tau_{i+1}-\tau_{i})^{2/3}$ per phase, where $c_{\rm{oracle}}>0$ is a numerical constant, yielding
@@
However, this oracle relies on complete knowledge of $\tau_{i}$’s. In contrast, our algorithm MDBE is fully adaptive. Theorem 2 proves that it achieves minimax optimal regret, up to poly-logarithmic factors.
@@
Since $\tilde{L}_{T}\leq S_{T}\leq L_{T}$, Theorem 2 achieves minimax optimal rates of $\tilde{\mathcal{O}}(S_{T}^{1/3}T^{2/3})$ and $\tilde{\mathcal{O}}(L_{T}^{1/3}T^{2/3})$, without requiring prior knowledge of $S_{T}$ or $L_{T}$, respectively. The bound provided by Theorem 2 can be significantly sharper when $\tilde{L}_{T}\ll S_{T}$. The following result further shows that Theorem 2 also recovers the optimal rate in terms of the total variation $V_{T}$. The proof is deferred to Appendix D.
@@
5 Proof intuition of Theorem 2
@@
A natural question is whether one could directly apply the MASTER procedure of Wei and Luo (2021) using a discretization of the arm space as the base algorithm. However, such a direct application would not be possible in our setting, for the following reasons. The MASTER procedure is specifically designed to detect and adapt to changes in the mean reward functions $\mu_{t}$, whereas our goal is to handle significant shifts, defined in terms of cumulative reward gaps $\sum_{t}\delta_{t}$. This distinction is crucial: the quantity $\sum_{t}\delta_{t}$ is inherently more stable than the instantaneous mean rewards, as it captures suboptimality over time, even when the underlying reward functions change smoothly or gradually. As a result, the detection problem in our setting is more subtle and requires finer control over the accumulation of regret. Therefore, even when discretizing the arm space with replays at different depths, the analysis of Wei and Luo (2021) would not achieve the desired dependence on $\tilde{L}_{T}$.
@@
We further emphasize that our analysis goes beyond a straightforward adaptation of the techniques developed for the $K$-armed setting by Suk and Kpotufe (2022). In their setting, the algorithm maintains a set of safe arms and applies a variant of Successive Elimination (Even-Dar et al., 2006) by uniformly sampling from the current set and progressively eliminating arms identified as suboptimal. Roughly speaking, to detect changes of magnitude $\Delta$, they schedule replays of duration $d=K/\Delta^{2}$, during which all $K$ arms are re-sampled and their estimates updated. However, such a replay scheme is infeasible in our infinite-arm setting. If we were to discretize the arm space into $2^{m}$ bins at each replay, reliably detecting a shift of magnitude $\Delta$ would require roughly $2^{m}/\Delta^{2}$ rounds, resulting in a prohibitive regret cost. To circumvent this issue, our algorithm schedules replays at different discretization scales. By leveraging the hierarchical structure of the dyadic tree $\mathcal{T}$, we update estimates of the gaps at multiple depths simultaneously. We exploit the key observation that each sample $x_{t}$ provides information not only for its bin at depth $d$, but also for all its child bins at depths $d^{\prime}\geq d$. Through systematic aggregation of information across these nested bins, our multi-scale estimation strategy allows the algorithm to detect significant shifts efficiently at different resolutions, while avoiding the prohibitive sample complexity associated with naive discretization.
@@
We begin by highlighting the technical challenges that arise when estimating these cumulative gaps, in particular how they differ from estimating changes in the mean reward itself (Section 5.1). We then present the high-level ideas for bounding the dynamic regret in two scenarios: (i) when the process is in a stable block, and no significant shift is detected, we show that the replays do not increases the regret excessively (Section 5.2); and (ii) when a significant shift does occur, we show that the algorithm responds appropriately (Section 5.3). The proofs are provided in Appendix D.
@@ 5.1 Estimating the reward gaps
The continuous arm space setting presents key technical challenges. These difficulties are specific to our framework and are not addressed in existing work on non-stationary bandits. We believe these challenges are of independent interest and may be relevant beyond our specific setting.
Bias in mean reward estimation. Our algorithm actively prunes bins at different depths, meaning that for a given bin $B\in\mathcal{T}_{d}$, its sub-bins $B^{\prime}\subset B$ at deeper depths $d^{\prime}>d$ may be evicted. As a result, the empirical mean reward $\hat{\mu}_{t}(B)$ can be biased, since arms $x_{t}$ are only drawn from non-evicted sub-bins. Leveraging Assumption 1, we control this bias, which decays as
@@
Concentration of mean reward estimates. Due to the martingale property of $(\hat{\delta}_{t}(B^{\prime},B))_{t}$, we can tightly control its deviation from its conditional expectation. Recall that our hierarchical sampling scheme (Algorithm 2) ensures that the algorithm selects an active bin at the minimum active depth $d_{0}(t)$ and recursively samples from child bins at finer active depths. This choice guarantees that
@@
Let $\mathcal{E}_{1}$ be the following event: for all intervals $[s_{1},s_{2}]$, depths $d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}$, and bins $B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, we have
@@
Unlike in classical finite-armed bandits (Suk and Kpotufe, 2022, 2023; Suk and Agarwal, 2023), uniform sampling only holds strictly at depth $d_{0}(t)$. This subtle distinction plays a critical role in the analysis. Thanks to both the bias control (3) and the concentration control (4), we can ensure that bins are only evicted when truly unsafe, despite the complexity introduced by the hierarchical sampling.
@@ 5.2 Regret within a block without significant shift
Assume, for the sake of clarity, that no significant shift occurs during the block $[\tau_{l,m},\tau_{l,m+1}[$. When only the replay at depth $m$ is active, i.e. $\mathcal{D}_{t}=\{m\}$, the algorithm effectively runs a Successive Elimination strategy at depth $m$: at each round, it selects a bin uniformly at random from the set of safe bins $\mathcal{B}_{\mathrm{MASTER}}(m)$, samples an arm $x_{t}$ from this bin, and eliminates bins that are deemed unsafe at this depth. In this regime, since the discretization level at depth $m$ is tuned to the block length $\tau_{l,m+1}-\tau_{l,m}$, classical results indicate that the regret incurred within the block is of order $(8^{m})^{2/3}=4^{m}$ (up to logarithmic factors).
It remains to control the regret from replays across different scales. A key difficulty is that such replays may overlap, making it delicate to control the cumulative regret contribution at each depth simultaneously. A crucial property leveraged in the analysis is that our importance-weighted estimator (2) allows each sampled arm $x_{t}$ to update the mean estimates $\hat{\mu}_{t}(B)$ for all bins $B$ containing $x_{t}$. This multi-scale update property ensures that replays at any given depth $d$ contribute information not only at that level but also propagate to finer depths $d^{\prime}>d$. In particular, we leverage Proposition 1, which guarantees that if a bin is identified as unsafe at any depth $d$, then all arms it contains are suboptimal.
@@ 5.3 Detecting significant shifts
When a significant shift $\tau_{i}$ occurs within a block $[\tau_{l,m},\tau_{l,m+1}[$, the primary challenge is that the shift is initially undetected, potentially causing the algorithm to suffer large regret if it continues exploiting outdated information. In particular, such regret can accumulate if a bin that was previously evicted (due to being deemed unsafe) becomes optimal after the shift, while the algorithm keeps ignoring it. More precisely, detecting a shift of magnitude $\Delta$ at depth $d$ requires a replay of length approximately $2^{d}/\Delta^{2}$. Undetected shifts of this magnitude are tolerable at this resolution over a period of duration $D$ as long as $\Delta D\leq 4^{d}$. By carefully calibrating the probability of scheduling replays at each depth (see (1)), we guarantee that replays at depth $d$ occur frequently enough to ensure, with high probability, the timely detection of shifts of size $\Delta$ before their contribution to regret becomes significant.
@@ 6 Conclusion
Our work is primarily theoretical and does not address the computational efficiency or empirical performance of the proposed algorithm. Bridging this gap remains an important direction for future work. In particular, developing computationally efficient, adaptive algorithms that do not rely on prior knowledge of the non-stationarity budget (Gerogiannis et al., 2025) remains an open challenge. Additionally, extending the framework of significant shift detection to other structured bandit settings, such as linear or convex bandits, is an exciting avenue for future research.

### v2

@@ 4 Main results
[…]
@@
To give intuition whether the optimal rate $\mathcal{O}(\tilde{L}^{1/3}T^{2/3})$ is achievable, consider an oracle algorithm $\pi_{\rm{oracle}}$ with access to the exact times of the significant shifts $\{\tau_{i}\}_{i=0}^{\tilde{L}_{T}}$. The optimal algorithm partitions the horizon into $\tilde{L}_{T}$ phases and, within each phase, runs a minimax-optimal bandit algorithm (e.g. UCB [Auer et al., 2002a]) using a discretization of $(\tau_{i+1}-\tau_{i})^{1/3}$ arms. This oracle algorithm suffers regret at most $c_{\rm{oracle}}(\tau_{i+1}-\tau_{i})^{2/3}$ per phase, where $c_{\rm{oracle}}>0$ is a numerical constant, yielding
@@
However, this oracle relies on complete knowledge of $\tau_{i}$’s. In contrast, our algorithm MDBE is fully adaptive. Theorem˜2 proves that it achieves minimax optimal regret, up to poly-logarithmic factors.
@@
Since $\tilde{L}_{T}\leq S_{T}\leq L_{T}$, Theorem˜2 achieves minimax optimal rates of $\tilde{\mathcal{O}}(S_{T}^{1/3}T^{2/3})$ and $\tilde{\mathcal{O}}(L_{T}^{1/3}T^{2/3})$, without requiring prior knowledge of $S_{T}$ or $L_{T}$, respectively. The bound provided by Theorem˜2 can be significantly sharper when $\tilde{L}_{T}\ll S_{T}$. The following result further shows that Theorem˜2 also recovers the optimal rate in terms of the total variation $V_{T}$. The proof is deferred to Appendix˜D.
Remark 1 (Beyond $1$-Lipschitz bandits).
Minimax optimality of Theorem˜2 extends beyond Lipschitz bandits with Lipschitz constant $1$. In particular, for $(\kappa,\beta)$-Hölder bandits we can show that the minimax optimal rate (up to polylogarithmic factors) can be adapted as $\mathbb{E}\left[R(\pi_{\texttt{MBDE}},T)\right]\leq\widetilde{\mathcal{O}}\Big(T^{\frac{1+\beta}{1+2\beta}}\tilde{L}_{T}^{\frac{\beta}{1+2\beta}}\kappa^{\frac{1}{1+2\beta}}\Big)$, where the minimax optimality is with respect to $T$, $\tilde{L}_{T}$ and Hölder constants $\kappa$ and $\beta)$ jointly. The necessary modifications to our setting and algorithm are detailed in Appendix˜G.
@@
5 Proof intuition of Theorem˜2
@@
A natural question is whether one could directly apply the MASTER procedure of Wei and Luo [2021] using a discretization of the arm space as the base algorithm. However, such a direct application would not be possible in our setting, for the following reasons. The MASTER procedure is specifically designed to detect and adapt to changes in the mean reward functions $\mu_{t}$, whereas our goal is to handle significant shifts, defined in terms of cumulative reward gaps $\sum_{t}\delta_{t}$. This distinction is crucial: the quantity $\sum_{t}\delta_{t}$ is inherently more stable than the instantaneous mean rewards, as it captures suboptimality over time, even when the underlying reward functions change smoothly or gradually. As a result, the detection problem in our setting is more subtle and requires finer control over the accumulation of regret. Therefore, even when discretizing the arm space with replays at different depths, the analysis of Wei and Luo [2021] would not achieve the desired dependence on $\tilde{L}_{T}$.
@@
We further emphasize that our analysis goes beyond a straightforward adaptation of the techniques developed for the $K$-armed setting by Suk and Kpotufe [2022]. In their setting, the algorithm maintains a set of safe arms and applies a variant of Successive Elimination [Even-Dar et al., 2006] by uniformly sampling from the current set and progressively eliminating arms identified as suboptimal. Roughly speaking, to detect changes of magnitude $\Delta$, they schedule replays of duration $d=K/\Delta^{2}$, during which all $K$ arms are re-sampled and their estimates updated. However, such a replay scheme is infeasible in our infinite-arm setting. If we were to discretize the arm space into $2^{m}$ bins at each replay, reliably detecting a shift of magnitude $\Delta$ would require roughly $2^{m}/\Delta^{2}$ rounds, resulting in a prohibitive regret cost. To circumvent this issue, our algorithm schedules replays at different discretization scales. By leveraging the hierarchical structure of the dyadic tree $\mathcal{T}$, we update estimates of the gaps at multiple depths simultaneously. We exploit the key observation that each sample $x_{t}$ provides information not only for its bin at depth $d$, but also for all its child bins at depths $d^{\prime}\geq d$. Through systematic aggregation of information across these nested bins, our multi-scale estimation strategy allows the algorithm to detect significant shifts efficiently at different resolutions, while avoiding the prohibitive sample complexity associated with naive discretization.
@@
We begin by highlighting the technical challenges that arise when estimating these cumulative gaps, in particular how they differ from estimating changes in the mean reward itself (Section˜5.1). We then present the high-level ideas for bounding the dynamic regret in two scenarios: (i) when the process is in a stable block, and no significant shift is detected, we show that the replays do not increases the regret excessively (Section˜5.2); and (ii) when a significant shift does occur, we show that the algorithm responds appropriately (Section˜5.3). The proofs are provided in Appendix˜D.
@@ 5.1 Estimating the reward gaps
The continuous arm space setting presents key technical challenges. These difficulties are specific to our framework and are not addressed in existing work on non-stationary bandits, and we believe these challenges are of independent interest and may be relevant beyond our specific setting.
The main challenge in this setting lies in estimating a continuous, non-stationary mean reward function without prior knowledge of the magnitude of distributional shifts, or, equivalently, the discretization level and time scale at which such shifts become detectable. Coarse discretization is essential to rapidly detect large shifts, while finer discretization is required to capture smaller yet statistically significant changes. This motivates the multi-scale replay framework of MDBE in which each scale contributes adaptively based on the magnitude of the underlying shift.
Bias in mean reward estimation. Our algorithm actively prunes bins at different depths, meaning that for a given bin $B\in\mathcal{T}_{d}$, its sub-bins $B^{\prime}\subset B$ at deeper depths $d^{\prime}>d$ may be evicted. As a result, the empirical mean reward $\hat{\mu}_{t}(B)$ can be biased, since arms $x_{t}$ are only drawn from non-evicted sub-bins. Leveraging Assumption˜1, we control this bias, which decays as
@@
Concentration of mean reward estimates. Due to the martingale property of $(\hat{\delta}_{t}(B^{\prime},B))_{t}$, we can tightly control its deviation from its conditional expectation. Recall that our hierarchical sampling scheme (Algorithm˜2) ensures that the algorithm selects an active bin at the minimum active depth $d_{0}(t)$ and recursively samples from child bins at finer active depths. This choice guarantees that
@@
Let $\mathcal{E}_{1}$ be the following event: for all intervals $s_{1},s_{2}]$, depths $d\in\bigcap_{t=s_{1}}^{s_{2}}\mathcal{D}_{t}$, and bins $B,B^{\prime}\in\mathcal{B}_{[s_{1},s_{2}]}(d)$, we have
@@
Unlike in classical finite-armed bandits [Suk and Kpotufe, 2022, 2023, Suk and Agarwal, 2023], uniform sampling only holds strictly at depth $d_{0}(t)$. This subtle distinction plays a critical role in the analysis. Thanks to both the bias control (3) and the concentration control (4), we can ensure that bins are only evicted when truly unsafe, despite the complexity introduced by the hierarchical sampling.
@@ 5.2 Regret within a block without significant shift
Assume, for the sake of clarity, that no significant shift occurs during the block $\tau_{l,m},\tau_{l,m+1}[$. When only the replay at depth $m$ is active, i.e. $\mathcal{D}_{t}=\{m\}$, the algorithm effectively runs a Successive Elimination strategy at depth $m$: at each round, it selects a bin uniformly at random from the set of safe bins $\mathcal{B}_{\mathrm{MASTER}}(m)$, samples an arm $x_{t}$ from this bin, and eliminates bins that are deemed unsafe at this depth. In this regime, since the discretization level at depth $m$ is tuned to the block length $\tau_{l,m+1}-\tau_{l,m}$, classical results indicate that the regret incurred within the block is of order $(8^{m})^{2/3}=4^{m}$ (up to logarithmic factors).
It remains to control the regret from replays across different scales. A key difficulty is that such replays may overlap, making it delicate to control the cumulative regret contribution at each depth simultaneously. A crucial property leveraged in the analysis is that our importance-weighted estimator (2) allows each sampled arm $x_{t}$ to update the mean estimates $\hat{\mu}_{t}(B)$ for all bins $B$ containing $x_{t}$. This multi-scale update property ensures that replays at any given depth $d$ contribute information not only at that level but also propagate to finer depths $d^{\prime}>d$. In particular, we leverage Proposition˜1, which guarantees that if a bin is identified as unsafe at any depth $d$, then all arms it contains are suboptimal.
@@ 5.3 Detecting significant shifts
When a significant shift $\tau_{i}$ occurs within a block $\tau_{l,m},\tau_{l,m+1}[$, the primary challenge is that the shift is initially undetected, potentially causing the algorithm to suffer large regret if it continues exploiting outdated information. In particular, such regret can accumulate if a bin that was previously evicted (due to being deemed unsafe) becomes optimal after the shift, while the algorithm keeps ignoring it. More precisely, detecting a shift of magnitude $\Delta$ at depth $d$ requires a replay of length approximately $2^{d}/\Delta^{2}$. Undetected shifts of this magnitude are tolerable at this resolution over a period of duration $D$ as long as $\Delta D\leq 4^{d}$. By carefully calibrating the probability of scheduling replays at each depth (see (1)), we guarantee that replays at depth $d$ occur frequently enough to ensure, with high probability, the timely detection of shifts of size $\Delta$ before their contribution to regret becomes significant.
@@ 6 Conclusion
While our work is mainly theoretical, we also analyze the computational worst-case complexity of MDBE in Appendix˜H and show its empirical performance on a synthetic example in Appendix˜I. Developing computationally efficient algorithms with strong empirical performance that do not require prior knowledge of the non-stationarity [Gerogiannis et al., 2025a, b] remains an open problem. Future work could extend the significant shift detection framework to other structured bandit models such as convex or linear bandits.
Broader impacts
This work is mainly theoretical and contributes to the analysis of algorithms for sequential-decision making under uncertainty. Our generic setting and algorithms have broad potential use; practitioners will therefore need to specifically address possible social impacts with respect to the relevant application.
@@ Acknowledgements
The authors gratefully acknowledge Joe Suk for his insightful discussions and for sharing his code, which served as inspiration for our numerical experiments.